OpenStudy (anonymous):

Add and simplify: ( t+2)/3 - (t-4)/6

OpenStudy (anonymous):

\[\frac{ 2(t+2)-(t-4) }{ 6 }\]

OpenStudy (anonymous):

\[\frac{ 2t+4-t+4 }{ 6 }\]

OpenStudy (anonymous):

\[\frac{ t+8 }{ 6}\]

OpenStudy (anonymous):

the trick is just getting common denominators

OpenStudy (anonymous):

Neither of those answers match the given answers...:\ @Nurali @robz8

OpenStudy (anonymous):

what are the answers then?

OpenStudy (anonymous):

-2/3 , t-1/3 , or t/2

OpenStudy (anonymous):

did you type the question in correctly?

OpenStudy (anonymous):

you typed it in wrong, you said - instead of +

OpenStudy (anonymous):

\[\frac{ 2t+4+t-4 }{ 6 }\]

OpenStudy (anonymous):

Oh my mistake, sorry.

OpenStudy (anonymous):

\[\frac{ 3t }{6 }\]

OpenStudy (anonymous):

t/2 is your answer

OpenStudy (anonymous):

Alright, thanks so much!

OpenStudy (anonymous):

you're welcome

OpenStudy (pradius):

lol :) watch those parenthesis :)